# Byung-Kwan Lee KAIST

### v2

@@ 1 Introduction
[…]
Figure 3: Plotting the regressed relationships between (a) C2B and B2C for each object category, (b) the average of C2B & B2C and zero-shot GQA (Hudson and Manning, 2019) performance for each object category, (c) the average of C2B & B2C and zero-shot TextVQA (Singh et al., 2019) performance for each object category to visualize their correlations.
@@
By employing the Crayon Prompt, we create simple crayon instructions to enhance object-level image understanding. Additionally, we utilize the visual instruction tuning datasets (Liu et al., 2023c, b; Chen et al., 2023d) for zero-shot VL tasks. However, naively conducting visual instruction tuning may result in the forgetting of object-level image understanding acquired from crayon instructions. Hence, we propose a learning strategy called Dual QLoRA involving two QLoRA (Dettmers et al., 2023) modules. One module is trained for crayon instructions while the other module for visual instruction tuning datasets is frozen, and vice versa. This approach enables efficient fusion of crayon instructions and visual instruction tuning datasets while preserving the capabilities of both object-level image understanding and complex question answering. Pursuing parameter-efficient training, we employ quantized LoRA (QLoRA) instead of normal LoRA (Hu et al., 2021).
@@
  • To the best of our knowledge, we first reveal the intriguing property of current VLMs, wherein object-level image understanding is strongly correlated with zero-shot VL performance.
@@ 2 Research Backgrounds
CPT focuses solely on learning semantic and numbering queries in the Crayon Prompt and its MLP connector with the MS-COCO 2017 dataset Lin et al. (2014), aligning them with the backbone MLM to enhance object-level image understanding of [Uncaptioned image] CoLLaVO. On the other hand, Crayon Prompt-based Instruction Tuning (CIT) utilizes the visual instruction tuning datasets (Liu et al., 2023c, b; Chen et al., 2023d) as well as crayon instructions to handle complex question answering for VL tasks. It involves training the semantic and numbering queries and the MLP connector again, this time along with the backbone MLM of [Uncaptioned image] CoLLaVO.
When training the MLM with CIT, we introduce a learning strategy called Dual QLoRA, which manages object-level image understanding and complex VL performance, respectively, to effectively maintain both aspects. Figure 6 provides an overview of Dual QLoRA, where Image-CIT denotes using crayon instructions to bootstrap object-level image understanding and training only the first QLoRA module, while VL-CIT indicates using complex question-answer pairs from visual instruction tuning datasets to achieve zero-shot VL performance and training only the second QLoRA module. During CIT, we present an image in the form of Crayon Prompt to [Uncaptioned image] CoLLaVO, and randomly determine whether to proceed with Image-CIT or VL-CIT. The overarching objective of Dual QLoRA is to efficiently preserve both capabilities of object-level image understanding and complex VL performance. Note that the key distinction between CPT and Image-CIT lies in whether the backbone MLM of [Uncaptioned image] CoLLaVO is trained or not. Further details on minor aspects will be addressed in the following section.
Figure 7: In (a) and (b), the mean accuracy over the Top-20 object categories is denoted by the blue triangles, that over the Bottom-20 is denoted by the green triangles, and the red squares represent the average of all categories to visualize object-level image understanding of VLMs. In (c), zero-shot performance of VLMs on several VL benchmarks (MME-P (1/20 of score), GQA, TextVQA, SQA-IMG, SEED-IMG (accuracy)) are shown.
@@ 4 Experiments
(a): we employ Quantized Low-Rank Adaptation (QLoRA) (Hu et al., 2021; Dettmers et al., 2023) since [Uncaptioned image] CoLLaVO pursues efficient training with minimal parameter tuning. Double quantization and normalized float 4-bit (nf4) are used with LoRA of $r=64$ and $\alpha=64$. (b): In contrast to CPT with only crayon instructions and images from MS-COCO 2017, CIT is conducted with visual instruction tuning datasets (Liu et al., 2023c, b; Chen et al., 2023d) as well. Hence, many images contain unrecognizable objects, such as text, code, posters, or mathematical symbols. Consequently, a panoptic color map with the unknown category and ‘0’ numbering will be generated, and the semantic query of the unk category and numbering query of ‘0’ will operate to create the Crayon Prompt in these cases. (c): Once the color map is given with discernible objects, text descriptions, including object names, their numbering indices, and their bounding box coordinates, are added to the question template. Conversely, if an image contains no objects, the question template includes the phrase “None of detailed object information for image.” (d): Regarding training, we train [Uncaptioned image] CoLLaVO with a batch size of 32 in one epoch using the AdamW (Loshchilov and Hutter, 2019) optimizer, scheduled by cosine annealing (Loshchilov and Hutter, 2016) from a learning rate of 1e-4 to 1e-6 for CPT and from 1e-5 to 1e-6 for CIT, respectively. (e): To find the best performance, [Uncaptioned image] CoLLaVO uses greedy or beam search ($n=3$) for text generation without any other hyper-parameters.
@@
Table 1: Evaluating zero-shot performances of [Uncaptioned image] CoLLaVO on eleven vision language datasets compared with the current powerful VLMs such as InstructBLIP, Qwen-VL, LLaVA1.5, and so forth.
@@
Before delving into validating [Uncaptioned image] CoLLaVO in VL tasks, it is crucial to ensure its proficiency in object-level image understanding. We assessed the accuracy of 80 object categories classified as ‘thing’ (See Appendix A) in the MS-COCO 2017 across two directions: Class2Binary (C2B) and Box2Class(B2C), using four strong baselines: BLIP2, InstructBLIP, Qwen-VL, and LLaVA1.5. As illustrated in Figure 7(a)-(b), [Uncaptioned image] CoLLaVO nearly outperforms the baselines in three cases: Top-20, Bot-20, and Avg for both C2B and B2C. Furthermore, it has the smallest performance gap between the Top-20 accuracy and the Bottom-20 accuracy for both C2B and B2C. Such observation indicates that [Uncaptioned image] CoLLaVO has a solid object-level image understanding across numerous object classes.
@@
Table 3: Controlling Dual QLoRA, Image-CIT, and VL-CIT in conducting CIT
@@
Following improved object-level image understanding, [Uncaptioned image] CoLLaVO is evaluated to measure zero-shot performance of VL tasks on renowned datasets (See Appendix B). As shown in Figure 1, 7(c), and Table 1, [Uncaptioned image] CoLLaVO surpasses several closed-source VLMs like GPT-4V, Gemini-Pro, Qwen-VL-Pro, as well as numerous open-source VLMs (See Appendix C for all VLMs used in evaluation). Particularly, noteworthy is its superiority over other models in the following benchmarks: MME, MM-Bench, MM-Bench-Chinese, and Q-Bench, which primarily evaluate visual perception and cognition abilities, where [Uncaptioned image] CoLLaVO demonstrates its significant advancements.
@@ 5 Discussion and Conclusion
We have shown the effectiveness of [Uncaptioned image] CoLLaVO alongside the Crayon Prompt and Dual QLoRA serving as a key in enhancing the object-level image understanding. Notably, Figure 8(a) illustrates the impressive ability of [Uncaptioned image] CoLLaVO achieving cutting-edge zero-shot performance with a relatively small size, thanks to its grasp of object-level understanding validated in SEED-IMG (Li et al., 2023b) with 9 types of questions on spatial understandings of images. Even from the perspective of hallucination, Figure 8(b) and Appendix D demonstrate that [Uncaptioned image] CoLLaVO reduces hallucination due to improved object-level image understanding, satisfactorily compared to both closed-source and open-source VLMs on POPE (Li et al., 2023d) and HallusionBench (Liu et al., 2023a). This suggests that while many researchers have dramatically scaled up their models and curated their own visual instruction tuning datasets, tackling object-level image understanding proves to be an effective strategy in designing VLMs.
@@ 6 Limitations
Crayon Prompts, relying on a panoptic color map, which is an external source beyond VLMs, may be constrained by the performance of the segmentation model and its encompassing number of object classes. Despite this, we have achieved commendable scores across all zero-shot tasks. It is expected for [Uncaptioned image] CoLLaVO to further improve once it incorporates a plethora of visual prompts obtained from diverse sources like robust object classification or image captioning models (Lee et al., 2020, 2022, 2023; Kim et al., 2023c), object-centric causally human-interpretable information (Kim et al., 2021, 2023b), open object detection (Zhang et al., 2023a), visual grounding (Liu et al., 2023d; Ren et al., 2024), interactive or unsupervised segmentation (Kirillov et al., 2023; Kim et al., 2023a), and optical characteristic recognition model (Bautista and Atienza, 2022).
@@ Appendix A COCO Classes for Panoptic Color Map
***: Object class that is not classified as ‘thing’ (countable) but ‘stuff’ (uncountable)
Appendix B Zero-shot Vision Language Datasets used in Evaluation
@@
Appendix C Vision Language Models used in Evaluation
@@
  • Monkey (Li et al., 2023e) increases input resolution within Large Multimodal Model (LMM) by dividing high-resolution images into smaller patches processed independently by static visual encoders. Avoiding extensive pre-training, it achieves resolutions up to 1344×896 without pretraining, improving identification of small objects and dense text.
@@
Appendix D Detail of POPE dataset for Hallucination

### v3

@@ 1 Introduction
However, it is yet uncharted whether the current leading VLMs truly possess a comprehensive understanding of fine-grained object information, and how this understanding influences their zero-shot performance in VL tasks related to each object. Hence, we delve into the analysis of object-level image understanding and zero-shot performance in VL tasks across different objects. To illustrate the behavior of object-level image understanding, we employ four strong baselines: BLIP2 (Li et al., 2023c), InstructBLIP (Dai et al., 2023), LLaVA1.5 (Liu et al., 2023b), and Qwen-VL (Bai et al., 2023). We pose two types of simple questions to gauge their object-level understanding such as: (1) ‘Is there any {object name} in this image?’ (Class2Binary: C2B), and (2) ‘Which object is in the specified bounding box [$x_{\text{min}}$, $y_{\text{min}}$, $x_{\text{max}}$, $y_{\text{max}}$]?’ (Box2Class: B2C). We then evaluate the accuracy of their responses for 80 object categories (See Section 4 for more details) while assessing their zero-shot performance on VL tasks across the same set of categories.
Figure 3: Plotting the regressed relationships between (a) C2B and B2C for each object category, (b) the average of C2B & B2C and zero-shot GQA (Hudson and Manning, 2019) performance for each object category, (c) the average of C2B & B2C and zero-shot TextVQA (Singh et al., 2019) performance for each object category to visualize their correlations. The light-colored areas indicate the vertical span with the probability of confidence interval 0.95.
@@
By employing the Crayon Prompt, we create simple crayon instructions to enhance object-level image understanding. Additionally, we utilize the visual instruction tuning datasets (Liu et al., 2023c, b; Chen et al., 2023d) for zero-shot VL tasks. However, conducting visual instruction tuning only may be not sure for the grasp of object-level image understanding. Hence, we propose a learning strategy called Dual QLoRA involving two QLoRA (Dettmers et al., 2023) modules. One module is trained for crayon instructions while the other module for visual instruction tuning datasets is frozen, and vice versa. This approach enables efficient fusion of crayon instructions and visual instruction tuning datasets while preserving the capabilities of both object-level image understanding and complex question answering. Pursuing parameter-efficient training, we employ quantized LoRA (QLoRA) instead of LoRA (Hu et al., 2021).
@@
  • To the best of our knowledge, we first reveal the intriguing property of current VLMs, wherein object-level image understanding is strongly correlated with zero-shot VL tasks.
@@ 2 Research Backgrounds
CPT focuses solely on learning semantic and numbering queries in the Crayon Prompt and its MLP connector with the MS-COCO 2017 dataset Lin et al. (2014), aligning them with the backbone MLM to enhance object-level image understanding of [Uncaptioned image] CoLLaVO. On the other hand, Crayon Prompt-based Instruction Tuning (CIT) utilizes the visual instruction tuning datasets (Liu et al., 2023c, b; Chen et al., 2023d) as well as crayon instructions to handle complex question answering for VL tasks. It involves training the semantic and numbering queries and the MLP connector again, along with the backbone MLM of [Uncaptioned image] CoLLaVO.
When training the MLM with CIT, we introduce a learning strategy called Dual QLoRA, which manages object-level image understanding and complex VL performance, respectively, to effectively maintain both aspects. Figure 6 provides an overview of Dual QLoRA, where Image-CIT denotes using crayon instructions to bootstrap object-level image understanding and training only the first QLoRA module, while VL-CIT indicates using complex question-answer pairs from visual instruction tuning datasets to achieve zero-shot VL performance and training only the second QLoRA module. During CIT, we present an image in the form of Crayon Prompt to [Uncaptioned image] CoLLaVO, and randomly determine whether to proceed with Image-CIT or VL-CIT. The overarching objective of Dual QLoRA is to efficiently preserve both capabilities of object-level image understanding and complex VL performance. Note that the key distinction between CPT and Image-CIT lies in whether the backbone MLM of [Uncaptioned image] CoLLaVO is trained or not. Further details will be addressed in the following section.
Figure 7: In (a) and (b), there are three metrics for the mean accuracy over Top-20 object categories, Bottom-20, and average of all categories to visualize object-level image understanding of VLMs. In (c), zero-shot performances of VLMs on MME-P (1/20 scaled down of score), SQA-IMG, TextVQA, and SEED-IMG (accuracy) are shown.
@@ 4 Experiments
(a): we employ Quantized Low-Rank Adaptation (QLoRA) (Hu et al., 2021; Dettmers et al., 2023) since [Uncaptioned image] CoLLaVO pursues efficient training with minimal parameter tuning. Double quantization and normalized float 4-bit (nf4) are used with LoRA of $r=64$ and $\alpha=64$. (b): In contrast to CPT with only crayon instructions and images from MS-COCO 2017, CIT is conducted with visual instruction tuning datasets (Liu et al., 2023c, b; Chen et al., 2023d) as well. Hence, many images contain unrecognizable objects, such as text, code, posters, or mathematical symbols. Consequently, a panoptic color map with the unknown category and ‘0’ numbering will be generated, and the semantic query of the unk category and numbering query of ‘0’ will operate to create the Crayon Prompt in these cases. (c): Once the color map is given with discernible objects, text descriptions, including object names, their numbering indices, and their bounding box coordinates, are added to the question template. Conversely, if an image contains no objects, the question template includes the phrase “None of detailed object information for image.” (d): Regarding training, we train [Uncaptioned image] CoLLaVO with a batch size of 32 in one epoch using the AdamW (Loshchilov and Hutter, 2019) optimizer, scheduled by cosine annealing (Loshchilov and Hutter, 2016) from a learning rate of 1e-4 to 1e-6 for CPT and from 1e-5 to 1e-6 for CIT, respectively. In addition, $h=35$, $w=35$, and $d=4096$ are used in Figure 5. (e): To find the best performance, [Uncaptioned image] CoLLaVO uses greedy or beam search ($n=3$) for text generation without any other hyper-parameters.
@@
Table 1: Evaluating zero-shot performances of [Uncaptioned image] CoLLaVO on ten vision language datasets compared with the current powerful VLMs such as InstructBLIP, Qwen-VL, LLaVA1.5, and so forth.
@@
Before delving into validating [Uncaptioned image] CoLLaVO in VL tasks, it is crucial to ensure its proficiency in object-level image understanding. We assessed the accuracy of 80 object categories classified as ‘thing’ (See Appendix A) in the MS-COCO 2017 across two directions: Class2Binary (C2B) and Box2Class(B2C), using four strong baselines: BLIP2, InstructBLIP, Qwen-VL, and LLaVA1.5. As illustrated in Figure 7(a)-(b), [Uncaptioned image] CoLLaVO nearly outperforms the baselines in three cases: Top-20, Bottom-20, and Average for both C2B and B2C. Furthermore, it has the smallest performance gap between the Top-20 accuracy and the Bottom-20 accuracy for both C2B and B2C. Such observation indicates that [Uncaptioned image] CoLLaVO has a solid object-level image understanding across numerous object classes. Beyond its ability, Appendix B shows zero-shot object grounding performance of [Uncaptioned image] CoLLaVO for strong generalization to grounding-level understanding.
@@
Table 3: Controlling Dual QLoRA, Image-CIT, and VL-CIT in conducting CIT.
@@
Following improved object-level image understanding, [Uncaptioned image] CoLLaVO is evaluated to measure zero-shot performance of VL tasks on renowned datasets (See Appendix C). As shown in Figure 1, 7(c), and Table 1, [Uncaptioned image] CoLLaVO surpasses several closed-source VLMs like GPT-4V, Gemini-Pro, Qwen-VL-Pro, as well as numerous open-source VLMs (See Appendix D for all VLMs used in evaluation). Particularly, noteworthy is its superiority over other models in the following benchmarks: MME, MM-Bench, MM-Bench-Chinese, and Q-Bench, which primarily evaluate visual perception and cognition abilities, where [Uncaptioned image] CoLLaVO demonstrates its significant margins.
@@ 5 Discussion and Conclusion
We have shown the effectiveness of [Uncaptioned image] CoLLaVO alongside Crayon Prompt and Dual QLoRA serving as a key in enhancing the object-level image understanding. Notably, Figure 8(a) illustrates the impressive ability of [Uncaptioned image] CoLLaVO achieving cutting-edge zero-shot performance with a relatively small size, thanks to its grasp of object-level understanding validated in SEED-IMG (Li et al., 2023b) with 9 types of questions on spatial understandings of images. Even from the perspective of hallucination, Figure 8(b) and Appendix E demonstrate that [Uncaptioned image] CoLLaVO reduces hallucination due to improved object-level image understanding, satisfactorily compared to both closed-source and open-source VLMs on POPE (Li et al., 2023d) and HallusionBench (Liu et al., 2023a). This suggests that while many researchers have dramatically scaled up their models and curated their own visual instruction tuning datasets, tackling object-level image understanding proves to be an effective strategy.
@@ 6 Limitations
Crayon Prompts, relying on a panoptic color map, which is an external source beyond VLMs, may be constrained by the performance of the segmentation model and its encompassing number of object classes. Despite this, we have achieved commendable scores across all zero-shot tasks. It is expected for [Uncaptioned image] CoLLaVO to further improve once it incorporates a plethora of visual prompts obtained from diverse sources like robust object classification or image captioning models (Lee et al., 2020, 2022, 2023; Kim et al., 2023c), object-centric causally human-interpretable information (Kim et al., 2021, 2023b), open object detection (Zhang et al., 2023a), visual grounding (Liu et al., 2023d; Ren et al., 2024), interactive or unsupervised segmentation (Kirillov et al., 2023; Kim et al., 2023a), and optical characteristic recognition model (Bautista and Atienza, 2022). Beyond its limitation, we believe our promising direction for crayon prompt-like visual cues surely makes VLMs further improve on image understanding for human-like AGI.
@@ Appendix A COCO Classes for Panoptic Color Map
banner blanket bridge cardboard
counter curtain door floor-wood
flower fruit gravel house
light mirror net pillow
platform playingfield railroad river
road roof sand sea
shelf snow stairs tent
towel wall-brick wall-stone wall-tile
wall-wood water window-blind window
tree fence ceiling sky
cabinet table floor pavement
mountain grass dirt paper
food building rock wall
rug
*: Object class that is not classified as ‘thing’ (countable) but ‘stuff’ (uncountable)
Appendix B Grounding-level Image Understanding
RefCOCO RefCOCO+ RefCOCOg
Model val testA testB val testA testB val test
Shikra-7B 87.01 90.61 80.24 81.60 87.36 72.12 82.27 82.19
Ferret-7B 87.49 91.35 82.45 80.78 87.38 73.14 83.93 84.76
VistaLLM-7B 88.10 91.50 83.00 82.90 89.80 74.80 83.60 84.40
Qwen-VL-7B 89.36 92.26 85.34 83.12 88.25 77.21 85.58 85.48
CogVLM-Grounding-17B 92.76 94.75 88.99 88.68 92.91 83.39 89.75 90.79
[…]
Table 4: Comparing object grounding performances of Shikra (Chen et al., 2023c), Ferret (You et al., 2024), VistalLLM (Pramanick et al., 2023), Qwen-VL (Bai et al., 2023), CogVLM-Grounding (Wang et al., 2023), and CoLLaVO on several object grounding benchmarks: RefCOCO, RefCOCO+, and RefCOCOg (Kazemzadeh et al., 2014). Even though CoLLaVO did not use object grounding dataset (RefCOCO) in training phase, CoLLaVO shows comparable zero-shot object grounding performances, compared with (no zero-shot) other models specifically targeting object grounding task trained with RefCOCO grounding dataset.
Appendix C Zero-shot Vision Language Datasets used in Evaluation
@@
Appendix D Vision Language Models used in Evaluation
@@
Appendix E Detail of POPE dataset for Hallucination